OpenStudy (anonymous):

how to simplify this equation? cos 70cos10 + sin70sin10

OpenStudy (anonymous):

\(\cos(A - B) = \cos(A) \cos(B) + \sin(A) \sin(B) \)

OpenStudy (anonymous):

i still can`t get the answer.the answer is 1/2..

OpenStudy (anonymous):

cos 70cos10 + sin70sin10 fits somewhere in the above formula

OpenStudy (anonymous):

Cos(70-10) = cos(60) = 1/2

OpenStudy (anonymous):

ok,understand already..thanks

OpenStudy (anonymous):

fantastick